Krista Dandurand was introduced and welcomed as a permanent member of the group. She will be working evenings at Kilmer beginning September 12.
Andy: Recommends that each unit/branch begin to consider if keeping track of time spent searching items would be beneficial. A test run will be performed at Alexander.
Matt: Has received Systems reports of items to be relocated to FOLIO collection, but haven't had the support staff available to proceed with the project. Working with Andy, Penny and Library Interiors to assess what materials will be needed to relocate 2/3 of the circulating collection to the West Room.
Stacy/Annex: Continues to receive transfers from the LSM weeding and Special Collections Room projects as well as items from the Math X collection.
Eddie/LSM: Continues to send items from Special Collections Room to Annex, all items transferred to SC/UA have already been sent. Progress on weeding project continues. Shelving of overflow has been completed. Cleaning and organizing the cage in basement continues.
Rob/Alex: Shelf reading of 2A is has almost been completed, and then 2B will be started.
Jeff/Alex: There appears to be a discrepancy in the number of books charged to CS-ALEX and the actual number of books physically held by Collection Services. A report will be run to help locate the all of the items on the list. Jeff and Andy reminded everyone to only handle repair and re-labelling problems for your unit's items. Items from other units should be forwarded to their home location with a note explaining the damage, not sent directly to Collection Services
Kristin/Kilmer: Getting acclimated with Kilmer and RUL and is eager to be of assistance for projects.
Paul/GovDoc: Completed a mini-shift of microfiche at Alex to accommodate arrival of latest volumes. Backlog of shelving at Alex has been completed. Has started a mini-shift at LSM to accommodate international and foreign documents being transferred to collection from the stacks.
Marty/LSM: Busy processing transfers of items from stacks to Gov. Docs. Working on Lost Assumed list.
Megan/Physics: Busy hiring and training new employees.
Brian/Alex: busy with start of semester and return of students.
Mary/Alex: Working to establish a centralized location for Water Emergency supplies. Labeling of Reference continues. A New Books section has been created in lobby, items will be on display for two weeks and then replaced with a new batch.
Shirley/Alex: Continues searching lst/pd, lst/rpl and lst/assumed lists and has had some success in finding items on the lists.
Nicole/Chang: Busy processing reserve materials. Phase 1 of Serials Inventory Project will be completed later this week.
Eugene/SMLR: Basement area, housing rarely used items has been transformed considerably. After the meeting the group was invited down to view the changes.
Andy distributed revised copies of the documents to the group to review. Additions and corrections were made and finalized versions will forwarded to Nicole and Marc Forster to post on NBL page.
Andy distributed an equipment checklist, so that each unit could inventory their supply. A list of any missing items should be forwarded to any by the end of the month, so that a supply order can be placed.
Penny made some suggestions for the purchase additional equipment that would be very useful during a water emergency. She will forward details of the items to Andy for consideration.
WEDNESDAY, OCTOBER 13, 2004
LIBRARIES ANNEX, BUSCH CAMPUS